what is the minimum score needed to make on the 5th test so that the mean test score is at least 85? The first 4 test were 87, 92, 76, and 89
Answer by nerdybill(5404)

Let x = score on 5th test
.
(87+92+76+89+x)/5=85
(87+92+76+89+x)/5=85
(344+x)/5 = 85
344+x = 425
x = 81